# From Alignment to Advancement: Bootstrapping Audio-Language Alignment with Synthetic Data

### v2

@@ IV-B Model Selection and Training Setups
Our framework integrates two key foundation models: Llama-3.1-8B-Instruct333huggingface.co/meta-llama/Llama-3.1-8B-Instruct[30] and the compact Whisper-small444huggingface.co/openai/whisper-small[37], which contains about 240 million parameters. To connect these components, followed by [40, 15, 17, 18, 19], we developed a specialized modality adapter, featuring a two-layer transformer decoder architecture. This adapter uses 64 specialized vectors to capture and process audio features from the encoder’s hidden states. In our implementation strategy, we maintain the original architecture of both LLaMA and Whisper unchanged. Instead, we focused our training efforts solely on the modality adapter, which introduces approximately 22 million trainable parameters. The system processes information by first converting text instructions through LLaMA’s embedding layer, then combining them with the processed audio features.
@@ IV-C Baseline Models and Evaluation Setups
In this study, we incorporate widely used, well-documented, and fully open-source models as baselines for comparison. These include Qwen-Audio-Chat [1], Qwen2-Audio-Instruct [2], SALMONN [3] (including 7B and 13B versions), LTU [4], and LTU-AS [5], as well as proprietary commercial models like Gemini-Pro-1.5 [36]. Since most ALLMs are instruction-tuned for open-ended responses [1, 2, 3, 4, 5], they generate answers based on user queries rather than selecting from predefined choices. To align with this design, we extract relevant answers from model outputs using a structured evaluation approach. All experiments adopt a consistent greedy decoding strategy with a maximum output length of 512.
@@
Audio Question Answering Benchmark. In ClothoAQA, which includes both binary and non-binary classification tasks, we report accuracy as the primary metric.
Audio Reasoning Benchmark. We report weighted F1 score for multi-class classification tasks. For Synonym-Hypernym Test, we also report weighted precision and recall. In addition, the original MMAU benchmark uses micro-averaged accuracy as its evaluation metric. The SAKURA test set follows a multiple-choice format, and we evaluate model performance using three key metrics. We report overall accuracy across all questions, as well as separate accuracy scores for single-hop and multi-hop questions.
@@ IV-D Ablation Study
Second, to validate that data generated by the backbone LLM outperforms data generated by an external LLM under this setting, we also examine a scenario where a stronger LLM is used to generate synthetic data via our proposed pipeline, BALSa. Specifically, we select Gemini-1.5-Pro 555gemini-1.5-pro-latest to synthesize the positive-only, positive and negative, and combined training data. The same methodology is applied, but a more advanced model is used to generate the data.
@@ V-A Audio Question Answering Benchmark
This benchmark focuses on assessing a model’s fundamental understanding of audio. It evaluates capabilities such as sound event recognition, acoustic scene classification, temporal comprehension, and audio attribute identification. By designing diverse types of questions, this benchmark examines whether a model can accurately interpret various aspects of an audio signal and provide reasonable and correct answers based on the given auditory information. This allows us to gauge the model’s core competence in audio understanding. For this benchmark, we have collected and organized the following evaluation datasets.
Clotho-AQA: Clotho-AQA [47] is a widely used dataset designed for audio question answering (AQA) tasks, where models generate natural language answers based on audio signals and corresponding questions. The dataset is constructed through crowd-sourcing, with questions collected for each audio file and answers provided by multiple annotators. To ensure data reliability, only questions where all annotators unanimously agree on the answer are included. The final dataset consists of 789 instances.
Nonspeech7k AQA: Nonspeech7k [54] is a dataset of human non-speech sounds collected from real-life environments, covering seven sound categories: breath, cough, cry, laugh, scream, sneeze, and yawn. We utilize all 725 audio instances from the test split of the Nonspeech7k dataset for evaluation. To adapt the dataset to a question-answering (QA) format, we manually designed general questions such as: Classify the given sound as exactly one of: breath, cough, cry, laugh, scream, sneeze, or yawn. This ensures that the model follows explicit instructions when answering the questions.
CochlScene AQA: CochlScene [55] is a crowd-sourced dataset designed for acoustic scene classification, covering 13 different acoustic scenes. The primary goal of this task is to determine the environment in which a given audio recording was captured based on its acoustic characteristics. For evaluation, we sample 100 audio clips from each category in the test split, resulting in a total of 1,300 test instances. We converted the dataset into a question-answering (QA) format through the manual design of general questions like Which acoustic scene does this audio clip belong to? The answer must be one of the four given options: Residential Area, Restaurant, Subway Station, Car.
EDANSA-2019 AQA: The Ecoacoustic Dataset from Arctic North Slope Alaska (EDANSA-2019) [56] is an ecoacoustic dataset collected from the North Slope of Alaska during the summer of 2019. The dataset contains over 27 hours of labeled audio. Since this dataset was recorded in the Arctic, most existing models may not have been trained on similar data. Publicly available training datasets rarely include such recordings, making EDANSA-2019 an out-of-distribution (OOD) dataset. This poses a challenge for models, as they may struggle to generalize to these unfamiliar acoustic environments.
For evaluation, we selected sound events with a sufficient number of samples, covering fly, bird, woof, wind, rain, aircraft, and machine. Due to data limitations, fly (62 instances) and woof (15 instances) have fewer samples, while all other categories contain 200 randomly selected audio clips. In total, the test set consists of 1,077 instances. We reformulated the dataset into a question-answering (QA) format by manually crafting general questions, for example: Select the most appropriate category for the given sound from the following options: Wind, Machine, Rain, Fly.
@@ V-B Audio Reasoning Benchmark
This benchmark focuses on evaluating a model’s advanced reasoning capabilities in audio understanding. Beyond recognizing and interpreting audio content, models must demonstrate the ability to reason based on their understanding to make logical inferences. For this benchmark, we utilize the following dataset for this benchmark:
@@
Synonym and Hypernym Test: A previous study [57] introduced a synonym and hypernym test to assess whether models can understand audio and its semantic connections to text. Synonyms refer to words with similar meanings, while hypernyms denote broader categories that encompass more specific terms. These relationships enable the construction of text-based prompts that test a model’s understanding of similarity and hierarchy in audio reasoning. For example, given an audio clip of a songbird chirping, a possible question could be: Is the sound from an object that is a type of chordate? The expected answer is yes.
This benchmark evaluates whether audio-aware large language models genuinely comprehend audio content and semantic relationships, or if they merely rely on statistical associations between audio, keywords, and captions. We reproduce the evaluation setup and data selection from the original study, sourcing data from the AudioCaps dataset, resulting in 636 evaluation instances. To further analyze the model’s behavior, we separately compute weighted precision, recall, and F1 scores for questions where the correct answer is yes and no.
Audio Entailment: A previous study [58] introduced the task of audio entailment to assess a model’s ability to perform deductive reasoning. This task evaluates whether a textual description (hypothesis) can be logically inferred from an audio recording (premise). The conclusion falls into one of three categories: entailment, neutral, or contradiction, depending on the available evidence. Entailment occurs when the audio recording provides sufficient evidence to affirm the truth of the hypothesis. Neutral applies when the audio recording does not contain enough information to either confirm or deny the hypothesis. Contradiction is determined when the audio recording offers clear evidence that the hypothesis is false.
@@
We format this task as a multiple-choice question, requiring the model to identify the correct conclusion type based on the given prompt. For example, given an audio clip, the model is asked to select the hypothesis that corresponds to entailment: Based on what you hear, which of the following hypotheses is entailment? (A) A quiet room with no movement or wind. (B) A motorcycle is overtaking cars on a windy road. (C) Vehicles are moving on a road with wind noise present. In this setup, the order of the answer choices is randomly shuffled for each instance. The dataset sources audio recordings from Clotho [47] and AudioCaps [42]. For evaluation, we report the overall accuracy.
@@
SAKURA: This benchmark [50] evaluates multi-hop reasoning in audio understanding. We use the SAKURA test set from the audio track for this evaluation. In this benchmark, each audio clip is paired with two types of questions: single-hop and multi-hop. Single-hop questions directly ask about attributes of the sound event in the audio, such as Out of the animals listed below, can you identify the one that most closely matches the sound in the audio? In contrast, multi-hop questions require multiple steps of reasoning, combining fundamental audio understanding with text-based inference to arrive at the correct answer. For example, a multi-hop question might be: Taking into account the sound from the recording and the natural tendencies of this type of animal making the sound, which of the following behaviors is most representative of how it typically acts? (a) Wagging its tail (b) Jumping between branches (c) Hiding in shells (d) Flapping its wings.
@@
MMAU: This is a massive multi-task audio understanding and reasoning benchmark [51] designed to evaluate multimodal audio understanding models on tasks requiring expert-level knowledge and complex reasoning. In this study, we focus on assessing model performance specifically along the audio dimension, using the test-mini split for evaluation. The evaluation settings follow the original benchmark’s guidelines.
@@ V-C Reliability and Safety Benchmark
The Reliable and Safety Benchmark focuses on evaluating the trustworthiness and safety of audio-aware large language models. We select the following evaluation benchmark.
Audio Hallucination: Our previous work [13] introduced benchmarks to evaluate object hallucination in audio—a phenomenon where models incorrectly detect or infer non-existent sound events. In this task, given an audio clip, the model is tested on whether a specific sound event is present. Inspired by Polling-based Object Probing Evaluation (POPE) [59], our prior work adopted a similar methodology to evaluate ALLMs. We formulate object hallucination detection as a binary classification task, prompting ALLMs to respond with “Yes” or “No”. To assess model robustness, we design four different prompts that ask whether a particular sound is present, including variations such as Is there a sound of [object]?, Does the audio contain the sound of [object]?, Have you noticed the sound of [object]?, and Can you hear the sound of [object]?
This benchmark is constructed using AudioCaps, sampling audio clips that contain more than three sound events. It consists of 407 distinct audio clips, where we treat all ground truth labels as positive samples. To introduce negative samples, we apply a random negative sampling strategy, selecting three objects that are not present in the given audio. In addition, in order to further analyze the model’s behavior, we separately compute precision, recall, and F1 scores for questions where the correct answer is yes and no.
Instruction Following Ability Evaluation: Previous work [12] has assessed instruction-following capabilities and quantified catastrophic forgetting in audio-aware large language models. The findings highlight that most audio-aware large language models struggle with even basic instruction following, exhibiting significant performance degradation compared to text-based large language models. To evaluate the model’s instruction-following ability, we utilize both close-ended and open-ended questions. Close-ended questions require the model to adhere to specific formatting constraints, ensuring strict compliance with instructions. In contrast, open-ended questions allow for more flexibility, evaluating the model’s ability to follow complex and creative instructions. We report the instruction following rate (IFrate) and forgetting rate ($\Delta$), which is the relative IFrate difference between ALLMs and their LLM counterpart. A smaller forgetting rate indicates a lower degree of forgetting in ALLMs. Note that since we cannot directly access proprietary models like Gemini, we are unable to calculate the forgetting rate. Similarly, for the Qwen-Audio models [1, 2], which are finetuned from the pretrained Qwen-Audio base model [1], it is not feasible to compute the IFrate difference between ALLMs and their text-only LLM.
@@ VI-A Performance of BALSa
EDANSA ClothoAQA CochlScene NonSpeech Audio Hallucination
Models F1 Acc F1 F1 F1 (Y) F1 (N) F1 (W) Yes
Qwen-Audio-Chat[1] 34.50 74.90 55.58 34.73 75.25 52.97 64.11 81.04
Qwen2-Audio-Instruct[2] 41.40 75.55 60.16 54.84 72.89 54.44 63.67 75.39
SALMONN-7B[3] 26.36 76.56 37.99 78.29 70.06 26.48 48.27 92.14
SALMONN-13B[3] 29.96 83.99 31.06 77.24 76.94 60.07 68.50 76.78
LTU[4] 38.71 67.10 14.91 5.79 78.27 73.65 75.96 59.97
LTU-AS[5] 23.16 56.77 20.30 8.06 51.68 45.49 48.59 56.02
Gemini-1.5-Pro[36] 17.46 68.43 36.18 47.01 65.86 70.95 68.41 42.02
Ours (BALSa, only-positive) 45.57 75.11 39.28 54.17 73.05 55.15 64.10 74.81
Ours (BALSa, positive-negative) 42.06 74.96 42.87 53.66 80.18 74.04 77.1 63.69
Ours (BALSa, combined) 48.56 80.93 42.48 52.17 77.21 71.23 74.22 61.58
Ours (BALSa-MA, discrimination) 50.97 77.87 44.77 68.65 85.35 84.66 85.00 52.29
Ours (BALSa-MA, captioning) 53.80 78.75 44.47 70.03 84.00 84.93 84.47 47.01
In Table IV, compared to other publicly available baseline models, our proposed audio-aware large language models trained using the BALSa method (combined, positive-only, and positive-negative) achieve performance comparable to other public baseline models in audio question answering benchmarks. For instance, in ClothoAQA, our models perform on par with Qwen2-Audio-Instruct and SALMONN-7B. In NonSpeech AQA, our models attain results comparable to Qwen2-Audio-Instruct and Qwen-Audio-Chat. In CochlScene AQA, our models outperform SALMONN-7B and SALMONN-13B. Compared to Gemini-1.5-Pro, a proprietary model, our method remains competitive, leading by 6% to 25% on these evaluation benchmarks. While SALMONN-13B achieves the best performance on ClothoAQA and NonSpeech AQA, our models surpass it in other evaluation benchmarks. Despite falling behind Qwen-Audio-Chat and Qwen2-Audio-Instruct in tasks such as CochlScene AQA, our models demonstrate well-rounded performance across various benchmarks. These results highlight the effectiveness of the BALSa framework. By leveraging synthetic data to enhance audio-language alignment, BALSa delivers competitive results on traditional audio question answering benchmarks, often surpassing other baselines while maintaining consistent performance across different benchmarks. Notably, our models achieve these results using only 12% of the equivalent training duration required by the widely adopted and powerful Qwen2-Audio-Instruct model, as shown in Table III.
@@
The EDANSA AQA tasks serve as zero-shot benchmarks, evaluating models on out-of-domain audio that is not encountered during training. Our models, along with SALMONN and LTU-AS, are assessed under these conditions. Compared to other baselines, our method demonstrates a significant performance advantage. The results from this benchmark suggest that our approach exhibits a certain level of adaptability when applied to out-of-distribution data. Note that the Qwen series does not disclose its training datasets, thereby making it impossible to determine whether its results are obtained under a true zero-shot setting.
@@
In audio hallucination tasks, incorporating negative samples through the BALSa framework proves to be beneficial. After adding negative samples, the positive-negative setting improves the weighted F1 score by 13%, representing an approximate 20% increase compared to the positive-only setting. This suggests that a contrastive-like approach effectively help mitigate audio hallucinations while preserving performance on audio question answering. Notably, in audio hallucination tasks, many baseline models struggle to distinguish nonexistent sounds when the ground truth is no, frequently misidentifying them as present. In fact, most baseline models tend to answer yes regardless of whether the sound exists. This issue raises concerns about model reliability in real-world applications. Even models that perform well on other evaluation benchmarks, such as Qwen2-Audio-Instruct and Qwen-Audio-Chat, exhibit weaker performance in hallucination detection. Our proposed method leverages BALSa to generate negative samples. The addition of these samples, while simple, directly and effectively enhances performance on the audio hallucination benchmark. At the same time, it maintains performance on other evaluation benchmarks and, in some cases, even leads to slight improvements. This demonstrates that our approach not only mitigates hallucinations but also achieves a strong level of performance in fundamental audio understanding. Regarding the incorporation of negative samples, our experimental results indicate that separating positive and negative samples yields better performance compared to combining them into a single sample.

### v3

@@ IV-B Model Selection and Training Setups
Our framework integrates two key foundation models: LLaMA-3.1-8B-Instruct333huggingface.co/meta-llama/Llama-3.1-8B-Instruct[15] and the compact Whisper-small444huggingface.co/openai/whisper-small[52], which contains about 240 million parameters. In addition, to examine the generalizability of our approach across different language model backbones, we further trained a set of models by replacing Llama-3.1-8B-Instruct with Qwen3-4B 555huggingface.co/Qwen/Qwen3-4B [67], using several representative training data configurations. To connect these components, followed by [71, 59, 60, 29], we developed a specialized modality adapter, featuring a two-layer transformer decoder architecture. This adapter uses 64 specialized vectors to capture and process audio features from the encoder’s hidden states. In our implementation strategy, we maintain the original architecture of both LLaMA (or Qwen3-4B in the additional experiments) and Whisper unchanged. Instead, we focused our training efforts solely on the modality adapter, which introduces approximately 22 million trainable parameters. The system processes information by first converting text instructions through LLaMA’s embedding layer, then combining them with the processed audio features.
@@ IV-C Baseline Models and Evaluation Setups
In this study, we incorporate widely used, well-documented, and fully open-source models as baselines for comparison. These include Qwen-Audio-Chat [5], Qwen2-Audio-Instruct [6], Qwen2.5-Omni-Instruct [65], SALMONN [56] (including 7B and 13B versions), LTU [26], LTU-AS [25], GAMA [21], Audio Flamingo 2 [20], Audio Flamingo 3 [23], Audio-Reasoner [64], and Phi-4-Multimodal-Instruct [1], as well as proprietary commercial models like Gemini-Pro-1.5 [57], Gemini-Flash-2.5 [9], Gemini-Pro-2.5 [9], and GPT-4o-Audio [2]. Since most ALLMs are instruction-tuned for open-ended responses [5, 6, 56, 26, 25], they generate answers based on user queries rather than selecting from predefined choices. To align with this design, we extract relevant answers from model outputs using a structured evaluation approach. All experiments adopt a consistent greedy decoding strategy with a maximum output length of 512.
@@
Audio Question Answering Benchmark. Accuracy is adopted as the primary metric for ClothoAQA [42], which comprises both binary and non-binary classification tasks.
Audio Reasoning Benchmark. We report weighted F1 score for multi-class classification tasks. For Synonym-Hypernym Test, we also report weighted precision and recall. In addition, the original MMAU [54] benchmark uses micro-averaged accuracy as its evaluation metric. The subsequent MMAR [46] benchmark focuses on more challenging deep reasoning audio tasks, while adopting a similar evaluation approach to MMAU. The SAKURA [68] test set follows a multiple-choice format, and we evaluate model performance using three key metrics. We report overall accuracy across all questions, as well as separate accuracy scores for single-hop and multi-hop questions.
@@ IV-D Ablation Study
Second, to validate that data generated by the backbone LLM outperforms data generated by an external LLM under this setting, we also examine a scenario where a stronger LLM is used to generate synthetic data via our proposed pipeline, BALSa. Specifically, we select Gemini-1.5-Pro to synthesize the positive-only, positive and negative, and combined training data. The same methodology is applied, but a more advanced model is used to generate the data.
@@
In addition to the primary experiments, we conducted four complementary studies to evaluate different aspects of our framework: (1) replacing the LLaMA-3.1-8B backbone with Qwen3-4B to assess backbone generalizability; (2) applying LoRA fine-tuning instead of keeping the backbone frozen to study the effect of fine-tuning strategy; (3) varying prompts used in synthetic data generation to examine the impact of prompt design on model performance; and (4) data scaling analysis, where we evaluate performance under different proportions of the training set (10%, 25%, 50%, 75%, and 100%) to analyze the scalability of the BALSa approach.
@@ V-A Audio Question Answering Benchmark
This benchmark focuses on evaluating a model’s fundamental understanding of audio. It measures abilities like sound event recognition, acoustic scene classification, and temporal comprehension. By designing diverse instruction-following questions (e.g., Select the correct category for the sound: Wind, Machine, Rain, Fly.), the benchmark tests whether a model can accurately interpret different aspects of an audio signal and provide reasonable, correct answers based on the given auditory information. For this benchmark, we have collected and organized the following evaluation datasets.
Clotho-AQA: Clotho-AQA [42] is a widely used dataset designed for audio question answering (AQA) tasks, where models generate natural language answers based on audio and corresponding questions.
Nonspeech7k AQA: Nonspeech7k [53] is a dataset of human non-speech sounds collected from real-life environments, covering seven sound categories: breath, cough, cry, laugh, scream, sneeze, and yawn. We utilize all 725 audio instances from the test split of the Nonspeech7k dataset for evaluation.
CochlScene AQA: CochlScene [33] is a crowd-sourced dataset designed for acoustic scene classification, covering 13 different acoustic scenes. The primary goal of this task is to determine the environment in which a given audio recording was captured based on its acoustic characteristics. For evaluation, we sample 100 audio clips from each category in the test split, resulting in a total of 1,300 test instances.
EDANSA-2019 AQA: The Ecoacoustic Dataset from Arctic North Slope Alaska (EDANSA-2019) [7] contains recordings collected from the North Slope of Alaska in the summer of 2019. As an Arctic dataset, it represents an out-of-distribution domain rarely covered by public training corpora, posing a challenge for models to generalize to unfamiliar acoustic environments. For evaluation, we used sound events with sufficient samples: fly, bird, woof, wind, rain, aircraft, and machine. Due to limited data, fly (62 instances) and woof (15 instances) have fewer samples, while the remaining categories each include 200 randomly selected clips, totaling 1,077 instances.
@@ V-B Audio Reasoning Benchmark
This benchmark focuses on evaluating a model’s advanced reasoning capabilities in audio understanding. Models must go beyond recognition to reason and infer from audio. For this benchmark, we utilize the following datasets.
@@
Synonym and Hypernym Test: A previous study [8] introduced this test to examine whether models understand audio and its semantic ties to text. Synonyms capture similar meanings, while hypernyms capture broader categories. We leverage these relationships to construct prompts that probe similarity and hierarchy in audio reasoning. For example, given an audio of a songbird chirping, one question is: Is the sound from an object that is a type of chordate? (Answer: yes). This benchmark tests whether models truly capture semantic relationships in audio, rather than relying on surface statistical associations. Following the original setup with AudioCaps, we derive 636 evaluation instances and report weighted precision, recall, and F1 for yes/no questions.
@@
Audio Entailment: A previous study [12] introduced audio entailment to assess a model’s deductive reasoning: determining whether a textual hypothesis logically follows from an audio premise. Conclusions are labeled as entailment (audio affirms the hypothesis), contradiction (audio disproves it), or neutral (insufficient evidence). We frame this as a multiple-choice task where the model selects the correct conclusion. For example, given an audio clip, the prompt may ask: Which hypothesis is entailed? (A) A quiet room with no movement or wind. (B) A motorcycle is overtaking cars on a windy road. (C) Vehicles are moving on a road with wind noise present. Answer order is randomized. The dataset combines Clotho [42] and AudioCaps [35], and performance is measured by overall accuracy.
@@
SAKURA: The SAKURA benchmark [68] evaluates multi-hop reasoning in audio understanding, and we adopt its audio-track test set. Each audio clip is paired with two types of questions: single-hop and multi-hop. Single-hop questions ask directly about sound attributes, such as Out of the animals listed below, which most closely matches the audio? Multi-hop questions require combining audio understanding with text-based inference, e.g., Considering the sound and the natural tendencies of the animal, which behavior is most representative? (a) Wagging its tail (b) Jumping between branches (c) Hiding in shells (d) Flapping its wings.
@@
MMAU: We use the official MMAU-v05.15.25 benchmark [54], a massive multi-task audio understanding and reasoning benchmark designed to evaluate multimodal audio understanding models on tasks requiring expert-level knowledge and complex reasoning. In this study, we focus on assessing model performance specifically along the audio dimension, using the test-mini split for evaluation. The evaluation settings follow the original benchmark’s guidelines. 666github.com/Sakshi113/MMAU
MMAR: A large-scale multi-task audio reasoning benchmark [46] covering diverse real-world sound, music, and speech scenarios, collected from internet videos and refined through rigorous quality checks. It hierarchically categorizes questions into four reasoning layers (Signal, Perception, Semantic, Cultural), each requiring multi-step reasoning and, in some cases, domain-specific expertise. Evaluation follows the original benchmark setup.777github.com/ddlBoJack/MMAR
@@ V-C Reliability and Safety Benchmark
The Reliable and Safety Benchmark focuses on evaluating the trustworthiness and safety of ALLMs. We select the following evaluation benchmark.
Audio Hallucination: Our previous work [37] introduced benchmarks for object hallucination in audio, where models incorrectly infer non-existent sound events. Following the methodology of POPE [41], we cast hallucination detection as a binary task (Yes/No). To test robustness, we design four prompt variants (e.g., Is there a sound of [object]?, Does the audio contain…?). The benchmark is built from AudioCaps, sampling 407 clips with over three sound events; ground-truth objects are treated as positives, while three randomly chosen absent objects per clip serve as negatives. We report precision, recall, and F1 separately for yes and no cases.
Instruction Following Ability Evaluation: Prior work [43] shows that most ALLMs struggle with instruction following, exhibiting greater forgetting than text-only LLMs. We evaluate this ability using both close-ended (strict formatting) and open-ended (flexible, creative) questions. Performance is measured by instruction-following rate (IFrate) and forgetting rate ($\Delta$), where smaller $\Delta$ indicates less forgetting. Note that since we cannot directly access proprietary models like Gemini, we are unable to calculate the forgetting rate. Similarly, for the Qwen-Audio models [5, 6], which are finetuned from the pretrained Qwen-Audio base model [5], it is not feasible to compute the IFrate difference between ALLMs and their text-only LLM.
@@ VI-A Performance of BALSa
EDANSA ClothoAQA CochlScene NonSpeech Audio Hallucination Audio Entailment
Models F1 Acc F1 F1 F1 (Y) F1 (N) F1 (W) Yes Acc (C) Acc (A)
Qwen-Audio-Chat[5] 34.50 74.90 55.58 34.73 75.25 52.97 64.11 81.04 36.65 34.04
Qwen2-Audio-Instruct[6] 41.40 75.55 60.16 54.84 72.89 54.44 63.67 75.39 42.97 46.19
Qwen2.5-Omni-7B 70.81 86.90 81.24 54.03 89.72 85.81 87.77 49.59 56.81 62.81
SALMONN-7B[56] 26.36 76.56 37.99 78.29 70.06 26.48 48.27 92.14 33.24 34.10
SALMONN-13B[56] 29.96 83.99 31.06 77.24 76.94 60.07 68.50 76.78 35.25 30.81
LTU[26] 38.71 67.10 14.91 5.79 78.27 73.65 75.96 59.97 18.88 14.42
LTU-AS[25] 23.16 56.77 20.30 8.06 51.68 45.49 48.59 56.02 26.95 28.64
GAMA [21] 52.13 72.34 65.66 38.81 81.42 77.41 79.41 59.75 29.35 25.87
Audio Flamingo 2 [20] 28.65 83.41 38.45 10.62 68.35 20.43 44.39 93.08 58.41 55.31
Audio Flamingo 3 [23] 62.49 92.43 80.28 68.55 69.53 78.42 73.97 32.92 59.49 57.80
Audio-Reasoner [64] 57.95 70.89 61.18 83.75 71.61 73.51 72.59 36.12 10.02 23.54
Phi-4-Multimodal-Instruct [1] 32.20 66.38 31.49 9.66 70.53 33.38 51.95 88.66 50.08 49.38
Gemini-1.5-Pro[57] 17.46 68.43 36.18 47.01 65.86 70.95 68.41 42.02 23.32 22.87
Gemini-2.5-Flash [9] 48.31 71.62 68.80 65.64 83.08 81.21 82.14 49.75 17.94 16.27
Gemini-2.5-Pro [9] 62.79 82.18 57.89 77.60 86.35 68.81 77.58 44.90 17.84 18.27
GPT-4o-Audio [2] 64.47 87.36 75.73 74.07 84.62 50.71 67.67 51.8 18.50 22.80
BALSa (Ours)
– Llama-3.1, only-positive 45.57 75.11 39.28 54.17 73.05 55.15 64.10 74.81 50.88 52.41
– Llama-3.1, positive-negative 42.06 74.96 42.87 53.66 80.18 74.04 77.1 63.69 56.30 57.16
– Llama-3.1, combined 48.56 80.93 42.48 52.17 77.21 71.23 74.22 61.58 46.92 49.00
– Qwen3, only-positive 41.26 80.79 51.97 48.47 82.44 80.44 81.44 55.41 42.97 48.80
BALSa-MA (Ours)
– Llama-3.1, discrimination 50.97 77.87 44.77 68.65 85.35 84.66 85.00 52.29 54.67 55.66
– Llama-3.1, captioning 53.80 78.75 44.47 70.03 84.00 84.93 84.47 47.01 56.40 56.59
– Qwen3, discrimination 51.54 85.74 58.35 85.02 85.87 86.82 86.34 46.52 42.87 54.74
In Table IV, our proposed models trained using the BALSa method achieve competitive performance against a wide range of strong baseline models, especially when considering their training efficiency. For instance, in ClothoAQA, while models like Audio-Flamingo3 and Qwen-2.5-Omni set the state-of-the-art, our models perform robustly and are on par with other strong baselines like Qwen2-Audio. In benchmarks like NonSpeech and CochlScene AQA, although top-performing proprietary models like Gemini-2.5-Pro and models like Audio-Flamingo 3 lead, our models still attain solid results.
The primary advantage of the BALSa framework is its efficiency. While acknowledging the high performance of proprietary models like Gemini-2.5-Pro, GPT-4o, Audio Flamingo 3, and Qwen2.5-Omni it is crucial to highlight the training cost. As shown in Table III, our models achieve these competitive results using training data equivalent to only about 1,000 hours. This is a stark contrast to models like Qwen-2.5-Omni, which utilizes over 3.3 million hours of training data, and Audio Flamingo 3, which is trained on 54k hours. This highlights BALSa’s ability to achieve strong audio-language alignment with a fraction of the resources. In Figure 4, we compare average benchmark performance against training data size. Our framework demonstrates impressive data efficiency, achieving competitive accuracy under the smallest data scale among baseline models.
@@
The EDANSA AQA tasks serve as zero-shot benchmarks, evaluating models on out-of-domain audio that is not encountered during training. Our models, along with SALMONN and LTU-AS, are assessed under these conditions. Compared to other baselines evaluated under the same constraint, our method demonstrates a significant performance advantage. The results from this benchmark suggest that our approach exhibits a certain level of adaptability when applied to out-of-distribution data. Note that the training datasets for models like the Qwen series and various proprietary models are not fully disclosed, making it difficult to determine if their results are obtained under a true zero-shot setting.
@@
A standout strength of our method is its effectiveness in mitigating audio hallucinations. As shown in the Audio Hallucination benchmark, the BALSa framework, particularly with the inclusion of negative samples, proves highly beneficial. The positive-negative setting improves the weighted F1 score by 13%, an approximate 20% increase compared to the positive-only setting. Notably, even when compared against an expanded set of powerful models, our BALSa-MA variant (Qwen3-4B, discrimination) achieves one of the highest F1 (N) scores (86.82), outperforming strong models like Gemini-2.5-Pro (82.14). This is a critical advantage, as many baseline models that excel in other benchmarks, such as Qwen2-Audio-Instruct and Audio Flamingo 3, struggle to distinguish non-existent sounds. In fact, most baseline models tend to answer yes regardless of whether the sound exists. This issue raises concerns about model reliability in real-world applications. This result demonstrates that our approach not only mitigates hallucinations effectively but does so at a level competitive with or exceeding the state-of-the-art, highlighting its reliability for real-world applications while maintaining strong performance on other benchmarks.